OpenStudy (anonymous):

I need help solving the system of equations. 2x+y+z=10 2y+z=14 z+2

OpenStudy (blurbendy):

z+ 2 = ?

OpenStudy (anonymous):

i reckon she means z=2

OpenStudy (blurbendy):

i concur with your reckoning

OpenStudy (anonymous):

haha

OpenStudy (anonymous):

if z=2 2y+2=14 2y=12 y=6 2x+y+z=10 2x+6+2=10 2x=2 x=1

OpenStudy (anonymous):

Thank's!

OpenStudy (nurali):

2x+y+z=10---------(1) 2y+z=14-----------(2) z=2--------------(3) in equation '3' z=2 put 'z' in equation (2) 2y+2=14 subract both sides by '2' 2y+2-2=14-2 2y=12 y=6 put y and z in equation '1' 2x+y+z=10 2x+6+2=10 2x+8=10 2x=2 x=1

OpenStudy (anonymous):

Thank you!! I appreciate the help!

OpenStudy (nurali):

Anytime.